# Boundary Attention Constrained Zero-Shot Layout-To-Image Generation

### v1

@@ Abstract
Recent text-to-image diffusion models excel at generating high-resolution images from text but struggle with precise control over spatial composition and object counting. To address these challenges, several studies developed layout-to-image (L2I) approaches that incorporate layout instructions into text-to-image models. However, existing L2I methods typically require either fine-tuning pretrained parameters or training additional control modules for the diffusion models. In this work, we propose a novel zero-shot L2I approach, BACON (Boundary Attention Constrained generation), which eliminates the need for additional modules or fine-tuning. Specifically, we use text-visual cross-attention feature maps to quantify inconsistencies between the layout of the generated images and the provided instructions, and then compute loss functions to optimize latent features during the diffusion reverse process. To enhance spatial controllability and mitigate semantic failures in complex layout instructions, we leverage pixel-to-pixel correlations in the self-attention feature maps to align cross-attention maps and combine three loss functions constrained by boundary attention to update latent features. Comprehensive experimental results on both L2I and non-L2I pretrained diffusion models demonstrate that our method outperforms existing zero-shot L2I techniuqes both quantitatively and qualitatively in terms of image composition on the DrawBench and HRS benchmarks.
@@ 1 Introduction
Figure 1: An illustration of zero-shot layout-to-image (L2I) generation using various diffusion models. Text prompts and layout information, where specific concepts are localized within corresponding bounding boxes, are provided as input to the pretrained diffusion models to generate images that align with the instructions. Our method, Bacon, is capable of guiding non-L2I diffusion models such as SD [20] and SDXL [17] to generate images based on layout instructions, while also enhancing L2I models such as Gligen [11] to achieve improved spatial control.
@@
To tackle this deficiency, numerous studies [29, 3, 11, 32, 30] have explored layout-to-image (L2I) generation approaches that enable localization of concept positions in prompts using various forms of layout instructions such as semantic masks, bounding boxes, or sketches. By incorporating additional layout information, text-to-image (T2I) models can achieve more precise spatial controllability and generate datasets with specific ground-truth labels for data augmentation in supervised training [11, 29]. However, these L2I approaches require adapting the pretrained T2I models with additional control modules that need to be fine-tuned on large datasets containing paired images and layout annotations such as COCO [12].
@@
The significant data and computational requirements for model fine-tuning restrict the use of such approaches in many data-scarce or resource-constrained scenarios. To this end, a line of methods for layout-to-image generation in a zero-shot manner that require no additional supervised training has been proposed [31, 28, 27, 4, 16]. These techniques leverage the cross-attention maps extracted from U-Net [21] in the diffusion model to quantify the discrepancy between a synthetic image, sampled from the initialized latent feature $z_{t}$ based on the input text prompt, and the target layout instructions; the subsequent iterative update of the latent feature $z_{t}$ helps reduce the discrepancy. However, it has been observed that cross-attention maps typically exhibit high scores in the central regions of the concepts while assigning negligible scores to their edges [27, 31, 16]. Consequently, the generated concepts often appear larger than or misaligned with the specified bounding boxes, which diminishes the accuracy of layout control and leads to unsatisfactory image generation. Additionally, the existing approaches often face overlapping cross-attention maps when the bounding boxes for multiple objects within the same concept are closely spaced with minimal gaps, leading to inaccuracies in object counts relative to the input prompt.
@@
To improve spatial controllability and address semantic failures, we propose BACON, which stands for Boundary Attention Constrained generation, a novel zero-shot L2I approach. Specifically, we introduce two key design principles: (1) leveraging pixel-to-pixel correlations from visual self-attention maps to align the coarse-grained cross-attention maps, and (2) proposing a boundary attention constraint to address the previously discussed challenges of size misalignment and inaccurate counting. Self-attention maps capture pixel-to-pixel correlations in the visual features which can be used to filter noisy cross-attention maps and enhance the edges of concepts with low attention scores. Meanwhile, the boundary attention constraint ensures that cross-attention for each single object remains within the boundaries of the boxes, while promoting separation of the cross-attention maps of multiple objects within the same concept. Our comprehensive experimental results demonstrate that the proposed method outperforms existing zero-shot L2I methods both quantitatively and qualitatively in term of image compositions (spatial relationship, size, color, object counting) on the DrawBench [23] and HRS [1] benchmarks.
The main contribution of this paper can be summarized as follows:
@@
  • We investigate semantic failures in zero-shot L2I generation with complex layout inputs, focusing on the issue of overlapping cross-attention maps which leads to inaccurate object counting.
@@
  • We propose a novel method, BACON, that advances the L2I generation performance by incorporating self-attention enhancement to filter noisy cross-attention maps, and boundary attention constraints to prevent overlapping cross-attention maps.
@@
  • We perform comprehensive experiments comparing our method with the existing L2I approaches. The quantitative and qualitative experimental results demonstrate that our method achieves state-of-the-art performance as compared to the existing L2I techniques.
@@ 2.1 Text-To-Image Diffusion Models
Recently, diffusion models [7, 25, 26] have achieved remarkable success in image generation, demonstrating greater stability and controllability compared to GANs [5]. Essentially, a diffusion model learns the dynamics of mapping natural image-like data to a prior (e.g., Gaussian) distribution. The seminal work LDM [20] facilitates high-resolution image synthesis by performing forward and reverse processes in the latent space rather than the pixel space, thereby addressing issues of low inference speed and high training costs. In this framework, a U-net [21] is trained as the denoiser $\boldsymbol{\epsilon}_{\theta}$ to approximate the noise $\boldsymbol{\epsilon}_{t}$ based on the perturbed latent variable $\mathbf{z}_{t}$ and the condition $\mathbf{c}$ with the loss function
@@ 2.2 Layout-To-Image Generation
In addition to text prompts, layout-to-image (L2I) models require auxiliary layout instructions as input, i.e., semantic masks or bounding boxes. Several approaches aim to map layout instructions into the condition embedding space of diffusion models through supervised fine-tuning on datasets consisting of paired layout and image data. These approaches either integrate bounding boxes into the text prompts [29, 3, 32] and fine-tune the text encoder, or incorporate additional modules or adapters alongside the text encoder to enhance the understanding of layout instructions [11, 30]. Another line of zero-shot approaches aims to address the computational and data challenges associated with supervised L2I methods. Inspired by the connection between spatial layouts of generated images and the cross-attention maps observed in [6], DenseDiff [10] guides the placement of objects in targeted positions by manipulating the cross-attention maps of the text encoder. Follow-up studies such as A&E [2] and BoxDiff [28] propose optimizing the latent variable $z_{t}$ by maximizing the supremum of cross-attention scores located in specific regions during the reverse process of diffusion models. A&R [16] incorporates self-attention maps into the objective function to penalize objects that appear outside the designated boxes.
In contrast to previous studies that utilized the supremum, layout-guidance [32] uses the average of cross-attention scores in the objective function to update the latent variable $z_{t}$. Due to the typically coarse-grained and noisy nature of raw cross-attention maps, R&B [27] employs the Sobel Operator [8] to detect edges within these maps and select candidate boxes that encompass all edges for loss calculation. To address the additional computation and slower inference speed introduced by edge detection, concurrent work LoCo [31] directly leverages start-of-text tokens (SoT), end-of-text tokens (EoT), and self-attention maps to enhance cross-attention maps. However, none of the aforementioned approaches incorporate boundary attention constraints, leading to failures in object counting and precise spatial controllability.
@@ 3 Methodology
3.1 Zero-shot Layout-To-Image Generation
Zero-shot layout-to-image generation aims to generate images based on input text prompts and corresponding layout instructions using a pretrained diffusion model, without the need for additional parameter training or extra modules. Let us consider a text prompt $\mathbf{p}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{n}\}$ and a set of bounding boxes =$\mathcal{B}=\{\mathcal{B}_{i}\text{ for }\forall i\in\mathcal{I},\text{where }\mathcal{I}\subset[n]\}$; here $\mathcal{B}_{i}=\{\mathbf{b}^{(i)}_{1},\dots,\mathbf{b}^{(i)}_{N_{i}}\}$ denotes the corresponding bounding boxes for phrase $\mathbf{p}_{i}$ consisting of $N_{i}$ pairs of top-left and bottom-right points $(x_{1},y_{1},x_{2},y_{2})$ that specify locations of $N_{i}$ objects described by $\mathbf{p}_{i}$. The generated images should closely align with the text prompts while being consistent with the layout instructions defined by the set of bounding boxes $\mathcal{B}$. An illustration of zero-shot layout-to-image generation using bounding boxes as layout instruction is given in Figure 1.
Figure 2: The overall framework of BACON for zero-shot L2I generation consists of four steps: (a) aggregating cross-attention and self-attention maps obtained from the language-vision fusion blocks by layer-wise averaging; (b) enhancing cross-attention maps with self-attention maps; (c) computing three losses and combining them with coefficients; (d) optimizing $z_{t}$ by gradient descent in diffusion reverse process.
@@ 3.2 Cross-Attention and Self-Attention Maps
In the T2I diffusion models, text prompt $\mathbf{p}$ is encoded via text tokens $\mathbf{e}=f_{\text{CLIP}}(\mathbf{p})\in\mathbb{R}^{n\times d_{e}}$ using a pretrained CLIP encoder. These tokens are then input into the cross-attention layers where they are fused with the visual embedding characterized by $\mathbf{z}_{t}$. Specifically, in layer $l$, the visual and text embeddings are projected into the query $\mathbf{Q}_{z}^{l}\in\mathbb{R}^{hw\times d}$ and key $\mathbf{K}_{e}^{l}\in\mathbb{R}^{n\times d}$, respectively, allowing the cross-attention maps to be computed as
@@
which measure the pixel-to-pixel similarity in the visual features. These cross-attention and self-attention maps obtained from $L$ different cross-attention layers are aggregated as
@@
3.3 Boundary Attention Constrained Guidance
While prior studies [27, 4] demonstrated effectiveness of backward guidance, a procedure where cross-attention scores are used to form the loss utilized for optimizing the latent feature $z_{t}$ during the diffusion reverse process, the designed loss functions fail to address issues related to counting and semantic failures arising due to the coarse-grained nature of the cross-attention maps. In this section, we discuss the effect of attention enhancement and present three loss functions designed to address the problems encountered by previous schemes.
Figure 3: Visual comparison with different zero-shot L2I schemes based on Stable Diffusion XL [17]. The layout input and text prompt are shown in the first and second columns, respectively. The layout instructions are also annotated on the generated images with dashed boxes. Our method, BACON, significantly outperforms prior schemes in terms of spatial control and counting accuracy.
Attention enhancement. As depicted in the pink block in Figure 2, the raw cross-attention map for “cat” is coarse-grained, with attention dispersed across multiple parts of the object. It also shows notable scores for “dog” in the the cross-attention map for the “cat”, likely due to the high correlation between “cat” and “dog” within the pretrained model, which causes inconsistency in the loss computation. Inspired by the previous studies [14, 15] that aim to build synthetic datasets by generating grounded labels using cross-attention, we leverage the global information contained in self-attention maps to enhance the cross-attention maps,
@@ 3.3 Boundary Attention Constrained Guidance
Essentially, the self-attention map $\mathbf{A}_{s}$ captures the pairwise correlations between pixels, allowing salient parts in the raw cross-attention maps to propagate to the most relevant regions. Using these enhanced cross-attention maps, we compute three losses that are then utilized for improving the L2I generation.
@@
$\mathbf{M}^{(i)}[x,y]=\left\{\begin{aligned} &\;1,\text{ if }(x,y)\text{ inside }\mathcal{B}_{i}\\ &\;0,\text{ if }(x,y)\text{ outside }\mathcal{B}_{i}\\ \end{aligned}\right.,$ (7)
$L_{r}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{I}}\left(1-\frac{\mathbf{A}_{c}^{(i)}\odot\mathbf{M}^{(i)}}{\sum_{x,y}\mathbf{A}_{c}^{(i)}[x,y]}\right)^{2},$ (8)
where $\mathbf{A}_{c}^{(i)}[x,y]$ and $\mathbf{M}^{(i)}[x,y]$ are the (x,y)-th element in $\mathbf{A}_{c}^{(i)}$ and $\mathbf{M}^{(i)}$, respectively. However, relying solely on region-attention loss $L_{r}$ cannot ensure high-quality generation in the scenarios where multiple objects constrained by the same $\mathcal{B}_{i}$ have overlapping cross-attention scores, resulting in incorrect counting in the generated image.
Boundary-attention loss. As previously mentioned, adjacent bounding boxes for multiple objects described by the same phrase often lead to incorrect counting in the L2I generation due to the interference between cross-attention scores. However, the region-attention loss may remain small as long as these cross-attention scores are inside the target bounding boxes. We propose a boundary-attention loss that aims to isolate adjacent cross-attention maps corresponding to different objects,
$\mathbf{B}^{(i)}[x,y]=\left\{\begin{aligned} &\;1,\text{ if }(x,y)\text{ on }\mathcal{B}_{i}\\ &\;0,\text{ if }(x,y)\text{ not on }\mathcal{B}_{i}\\ \end{aligned}\right.,$ (9)
$L_{r}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{I}}\frac{\mathbf{A}_{c}^{(i)}\odot\mathbf{B}^{(i)}}{2(w_{\text{Box}}+h_{\text{Box}})},$ (10)
where $w_{\text{Box}}=\sum_{k=1}^{N_{i}}w_{k}^{(i)}$, $h_{\text{Box}}=\sum_{k=1}^{N_{i}}h_{k}^{(i)}$, and $w_{k}^{(i)}$ and $h_{k}^{(i)}$ denote the width and height of the bounding box $\mathbf{b}_{k}^{(i)}$, respectively. With the proposed boundary-attention loss, we are able to force $z_{t}$ to generate cross-attention maps without overlap between multiple objects and thus help improve the accuracy of counting.
Regularization loss. While the region-attention and boundary-attention losses generally help localize generated objects within the target bounding boxes, these losses may be minimal when an incorrect number of objects has cross-attention scores completely contained within the bounding boxes (i.e., experiencing no boundary crossing). To prevent bad generation in such scenarios, we propose a regularized loss that ensures the assigned objects do not have void cross-attention map,
@@
Latent Feature Optimization. At each sampling time step in the reverse process, $L_{bacon}$ can be computed as the weighted summation of $L_{\text{r}}$, $L_{\text{b}}$ and $L_{\text{reg}}$,
$L_{\text{bacon}}=L_{\text{r}}+\lambda L_{\text{b}}+\alpha L_{\text{reg}},$ (12)
where $\lambda$ and $\alpha$ control the intervention strength of $L_{b}$ and $L_{reg}$, respectively. We compute and use the gradient of $L_{\text{bacon}}$ to update $z_{t}$ as
$\hat{z_{t}}\xleftarrow{}z_{t}-\eta\nabla L_{\text{bacon}},$ (13)
where parameter $\eta$ controls the size of the updates. After $t_{\text{optim}}$ iterations or an early stop that happens when $L_{\text{bacon}}$ becomes smaller than a predetermined threshold, the optimized latent feature $\hat{z_{t}}$ is forwarded to the U-Net for predicting the latent feature $z_{t-1}$ at the previous time step.

### v2

# Training-Free Layout-to-Image Generation with Marginal Attention Constraints

@@ Abstract
Recently, many text-to-image diffusion models excel at generating high-resolution images from text but struggle with precise control over spatial composition and object counting. To address these challenges, prior works developed layout-to-image (L2I) approaches that incorporate layout instructions into text-to-image models. However, existing L2I methods typically require fine-tuning of pre-trained parameters or training additional control modules for the diffusion models. In this work, we propose a training-free L2I approach, MAC (Marginal Attention Constrained Generation), which eliminates the need for additional modules or fine-tuning. Specifically, we use text-visual cross-attention feature maps to quantify inconsistencies between the layout of the generated images and the provided instructions, and then compute loss functions to optimize latent features during the diffusion reverse process. To enhance spatial controllability and mitigate semantic failures in complex layout instructions, we leverage pixel-to-pixel correlations in the self-attention feature maps to align cross-attention maps and combine three loss functions constrained by boundary attention to update latent features. Comprehensive experimental results on both L2I and non-L2I pretrained diffusion models demonstrate that our method outperforms existing training-free L2I techniques both quantitatively and qualitatively in terms of image composition on the DrawBench and HRS benchmarks.
@@ 1 Introduction
Figure 1: An illustration of training-free layout-to-image (L2I) generation using various diffusion models. Text prompts and layout information, where specific concepts are localized within corresponding bounding boxes, are provided as input to the pretrained diffusion models to generate images that align with the instructions. Our method, MAC, is capable of guiding non-L2I diffusion models such as SD [22] and SDXL [19] to generate images based on layout instructions, while also enhancing L2I models such as Gligen [13] to achieve improved spatial control.
@@
To address this deficiency, numerous studies [31, 4, 13, 34, 32] have explored layout-to-image (L2I) generation approaches that enable the localization of concept positions in prompts using various forms of layout instructions such as semantic masks, bounding boxes, or sketches. By incorporating additional layout information, text-to-image (T2I) models can achieve more precise spatial controllability and generate datasets with specific ground-truth labels for data augmentation in supervised training [13, 31]. However, these L2I approaches require adapting the pretrained T2I models with additional control modules that must be fine-tuned on large datasets containing paired images and layout annotations such as COCO [14].
@@
The significant data and computational requirements for model fine-tuning restrict the use of such approaches in many data-scarce or resource-constrained scenarios. To this end, a series of methods for layout-to-image generation in a zero-shot manner, i.e., requiring no additional supervised training, have been proposed [33, 30, 29, 5, 18]. These techniques leverage the cross-attention maps extracted from U-Net [23] in the diffusion model to quantify the discrepancy between a synthetic image, sampled from the initialized latent feature $z_{t}$ based on the input text prompt, and the target layout instructions; the subsequent iterative update of the latent feature $z_{t}$ helps reduce the discrepancy. However, it has been observed that cross-attention maps typically exhibit high scores in the central regions of the concepts while assigning negligible scores to their edges [29, 33, 18]. Consequently, the generated concepts often appear larger than or misaligned with the specified bounding boxes, which diminishes the accuracy of layout control and leads to unsatisfactory image generation. Additionally, existing approaches often face overlapping cross-attention maps when the bounding boxes for multiple objects within the same concept are closely spaced with minimal gaps, leading to inaccuracies in object counts relative to the input prompt.
@@
To improve spatial controllability and address semantic failures, we propose MAC, which stands for Marginal Attention Constrained Generation, a novel training-free L2I approach. Specifically, we introduce two key design principles: (1) leveraging pixel-to-pixel correlations from visual self-attention maps to align the coarse-grained cross-attention maps, and (2) proposing a boundary attention constraint to address challenges of size misalignment and inaccurate counting. Self-attention maps capture pixel-to-pixel correlations in the visual features, which can be used to filter noisy cross-attention maps and enhance the edges of concepts with low attention scores. Meanwhile, the boundary attention constraint ensures that cross-attention for each single object remains within the boundaries of the boxes, while promoting the separation of the cross-attention maps of multiple objects within the same concept. Our comprehensive experimental results demonstrate that the proposed method outperforms existing training-free L2I methods both quantitatively and qualitatively in terms of image compositions (spatial relationship, size, color, object counting) on the DrawBench [24] and HRS [2] benchmarks. The main contribution of this paper can be summarized as follows:
@@
  • We investigate semantic failures in training-free L2I generation with complex layout inputs, focusing on the issue of overlapping cross-attention maps, which leads to inaccurate object counting.
@@
  • We propose a novel method, MAC, that advances the L2I generation performance by incorporating self-attention enhancement to filter noisy cross-attention maps, and boundary attention constraints to prevent overlapping cross-attention maps.
@@
  • We perform comprehensive experiments comparing our method with the existing L2I approaches. The quantitative and qualitative experimental results demonstrate that our method achieves state-of-the-art performance compared to existing L2I techniques.
@@ 2.1 Text-To-Image Diffusion Models
Recently, diffusion models [9, 26, 27] have achieved remarkable success in image generation, demonstrating greater stability and controllability compared to GANs [7]. Essentially, a diffusion model learns the dynamics of mapping natural image-like data to a prior (e.g., Gaussian) distribution. The seminal work LDM [22] facilitates high-resolution image synthesis by performing forward and reverse processes in the latent space rather than the pixel space, thereby addressing issues of low inference speed and high training costs. In this framework, a U-net [23] is trained as the denoiser $\boldsymbol{\epsilon}_{\theta}$ to approximate the noise $\boldsymbol{\epsilon}_{t}$ based on the perturbed latent variable $\mathbf{z}_{t}$ and the condition $\mathbf{c}$ with the loss
@@ 2.2 Layout-To-Image Generation
In addition to text prompts, layout-to-image (L2I) models require auxiliary layout instructions as input, i.e., semantic masks or bounding boxes. Several approaches aim to map layout instructions into the condition embedding space of diffusion models through supervised fine-tuning on datasets consisting of paired layout and image data. These approaches either integrate bounding boxes into the text prompts [31, 4, 34, 1] and fine-tune the text encoder, or incorporate additional modules or adapters alongside the text encoder to enhance the understanding of layout instructions [13, 32]. Another line of training-free approaches aims to address the computational and data challenges associated with supervised L2I methods. Inspired by the connection between spatial layouts of generated images and the cross-attention maps observed in [8], DenseDiff [12] guides the placement of objects in targeted positions by manipulating the cross-attention maps of the text encoder. Follow-up studies such as A&E [3] and BoxDiff [30] propose optimizing the latent variable $z_{t}$ by maximizing the supremum of cross-attention scores located in specific regions during the reverse process of diffusion models. A&R [18] incorporates self-attention maps into the objective function to penalize objects that appear outside the designated boxes.
In contrast to previous studies that utilized the supremum, layout-guidance [34] uses the average of cross-attention scores in the objective function to update the latent variable $z_{t}$. Due to the typically coarse-grained and noisy nature of raw cross-attention maps, R&B [29] employs the Sobel Operator [10] to detect edges within these maps and select candidate boxes that encompass all edges for loss calculation. To address the additional computation and slower inference speed introduced by edge detection, concurrent work LoCo [33] directly leverages start-of-text tokens (SoT), end-of-text tokens (EoT), and self-attention maps to enhance cross-attention maps. B2B [28] incorporates objection generation and attribute binding to improve the controllability. Concurrent work [6] explores the attention leakage problem and uses bounded attention for reducing semantic leakage. However, none of the aforementioned approaches incorporates marginal attention constraints, leading to failures in object counting and precise spatial controllability.
@@ 3 Methodology
3.1 Training-Free Layout-To-Image Generation
Training-free layout-to-image generation aims to generate images based on input text prompts and corresponding layout instructions using a pretrained diffusion model, without the need for additional parameter training or extra modules. Let us consider a text prompt $\mathbf{p}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{n}\}$ and a set of bounding boxes $\mathcal{B}=\{\mathcal{B}_{i}\text{ for }\forall i\in\mathcal{I},\text{where }\mathcal{I}\subset[n]\}$; here $\mathcal{B}_{i}=\{\mathbf{b}^{(i)}_{1},\dots,\mathbf{b}^{(i)}_{N_{i}}\}$ denotes the corresponding bounding boxes for phrase $\mathbf{p}_{i}$ consisting of $N_{i}$ pairs of top-left and bottom-right points $(x_{1},y_{1},x_{2},y_{2})$ that specify locations of $N_{i}$ objects described by $\mathbf{p}_{i}$. The generated images should closely align with the text prompts while being consistent with the layout instructions defined by the set of bounding boxes $\mathcal{B}$. An illustration of training-free layout-to-image generation using bounding boxes as layout instructions is given in Figure 1.
Figure 2: The overall framework of MAC for training-free L2I generation consists of four steps: (a) aggregating cross-attention and self-attention maps obtained from the language-vision fusion blocks by layer-wise averaging; (b) enhancing cross-attention maps with self-attention maps; (c) computing three losses and combining them with coefficients; (d) optimizing $z_{t}$ by gradient descent.
@@ 3.2 Cross-Attention and Self-Attention Maps
In the T2I diffusion models, text prompt $\mathbf{p}$ is encoded via text tokens $\mathbf{e}=f_{\text{CLIP}}(\mathbf{p})\in\mathbb{R}^{n\times d_{e}}$ using a pretrained CLIP encoder. These tokens are then input into the cross-attention layers, where they are fused with the visual embedding characterized by $\mathbf{z}_{t}$. Specifically, in layer $l$, the visual and text embeddings are projected into the query $\mathbf{Q}_{z}^{l}\in\mathbb{R}^{hw\times d}$ and key $\mathbf{K}_{e}^{l}\in\mathbb{R}^{n\times d}$, respectively, allowing the cross-attention maps to be computed as
@@
which measures the pixel-to-pixel similarity in the visual features. These cross-attention and self-attention maps obtained from $L$ different cross-attention layers are aggregated as
@@
3.3 Marginal Attention Constrained Guidance
While prior studies [29, 5] demonstrated effectiveness of backward guidance, a procedure where cross-attention scores are used to form the loss utilized for optimizing the latent feature $z_{t}$ during the diffusion reverse process, the designed loss functions fail to address issues related to counting and semantic failures arising due to the coarse-grained nature of the cross-attention maps. In this section, we discuss the effect of attention enhancement and present three loss functions to address the problems encountered by previous schemes.
Attention Enhancement. As depicted in the pink block in Figure 2, the raw cross-attention map for “cat” is coarse-grained, with attention dispersed across multiple parts of the object. It also shows notable scores for “dog” in the cross-attention map for the “cat”, likely due to the high correlation between “cat” and “dog” within the pretrained model, which causes inconsistency in the loss computation. Inspired by the previous studies [16, 17] that aim to build synthetic datasets by generating grounded labels using cross-attention, we leverage the global information contained in self-attention maps to enhance the cross-attention maps,
@@ 3.3 Marginal Attention Constrained Guidance
Essentially, the self-attention map $\mathbf{A}_{s}$ captures the pairwise correlations between pixels, allowing salient parts in the raw cross-attention maps to propagate to the most relevant regions. Using these enhanced cross-attention maps, we compute three losses for improving the L2I generation.
@@
[…]
$L_{\text{r}}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{I}}\left(1-\frac{\mathbf{A}_{c}^{(i)}\odot\mathbf{M}^{(i)}}{\sum_{x,y}\mathbf{A}_{c}^{(i)}[x,y]}\right)^{2},$ (8)
where $\mathbf{A}_{c}^{(i)}[x,y]$ and $\mathbf{M}^{(i)}[x,y]$ are the (x,y)-th element in $\mathbf{A}_{c}^{(i)}$ and $\mathbf{M}^{(i)}$, respectively. However, relying solely on region-attention loss $L_{\text{r}}$ cannot ensure high-quality generation in scenarios where multiple objects constrained by the same $\mathcal{B}_{i}$ have overlapping cross-attention scores, resulting in incorrect counting in the generated image.
Marginal-Attention Loss. As previously mentioned, adjacent bounding boxes for multiple objects described by the same phrase often lead to incorrect counting in the L2I generation due to the interference between cross-attention scores. However, the region-attention loss may remain small as long as these cross-attention scores are inside the target bounding boxes. We propose a marginal-attention loss that aims to isolate adjacent cross-attention maps corresponding to different objects,
[…]
$L_{\text{m}}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{I}}\frac{\mathbf{A}_{c}^{(i)}\odot\mathbf{B}^{(i)}}{2(w_{\text{Box}}+h_{\text{Box}})},$ (10)
where $w_{\text{Box}}=\sum_{k=1}^{N_{i}}w_{k}^{(i)}$, $h_{\text{Box}}=\sum_{k=1}^{N_{i}}h_{k}^{(i)}$, and $w_{k}^{(i)}$ and $h_{k}^{(i)}$ denote the width and height of the bounding box $\mathbf{b}_{k}^{(i)}$, respectively. With the proposed marginal-attention loss, we can force $z_{t}$ to generate cross-attention maps without overlap between multiple objects and thus help improve the accuracy of counting.
Regularization Loss. While the region-attention and marginal-attention losses generally help localize generated objects within the target bounding boxes, these losses may be minimal when an incorrect number of objects has cross-attention scores completely contained within the bounding boxes (i.e., experiencing no boundary crossing). To prevent bad generation in such scenarios, we propose a regularized loss that ensures the assigned objects do not have a void cross-attention map
@@
Latent Feature Optimization. At each sampling time step in the reverse process, $L_{\text{mac}}$ can be computed as the weighted summation of $L_{\text{r}}$, $L_{\text{m}}$ and $L_{\text{reg}}$,
$L_{\text{mac}}=L_{\text{r}}+\lambda L_{\text{m}}+\alpha L_{\text{reg}},$ (12)
where $\lambda$ and $\alpha$ control the intervention strength of $L_{\text{m}}$ and $L_{\text{reg}}$, respectively. We compute and use the gradient of $L_{\text{mac}}$ to update $z_{t}$ as $\hat{z_{t}}\xleftarrow{}z_{t}-\eta\nabla L_{\text{mac}},$ where parameter $\eta$ controls the size of the updates. After $t_{\text{optim}}$ iterations or an early stop that happens when $L_{\text{mac}}$ becomes smaller than a predetermined threshold, the optimized latent feature $\hat{z_{t}}$ is forwarded to the U-Net for predicting the latent feature $z_{t-1}$ at the previous time step.